# AI Agents vs. Agentic AI: A Conceptual Taxonomy, Applications and Challenges

### v3

@@ V Challenges and Limitations in AI Agents and Agentic AI
To systematically understand the operational and theoretical limitations of current intelligent systems, we present a comparative visual synthesis in Figure 12, which categorizes challenges and potential remedies across both AI Agents and Agentic AI paradigms. Figure 12a outlines the four most pressing limitations specific to AI Agents namely, lack of causal reasoning, inherited LLM constraints (e.g., hallucinations, shallow reasoning), incomplete agentic properties (e.g., autonomy, proactivity), and failures in long-horizon planning and recovery. These challenges often arise due to their reliance on stateless LLM prompts, limited memory, and heuristic reasoning loops.
In contrast, Figure 12b identifies eight critical bottlenecks unique to Agentic AI systems, such as inter-agent error cascades, coordination breakdowns, emergent instability, scalability limits, and explainability issues. These challenges stem from the complexity of orchestrating multiple agents across distributed tasks without standardized architectures, robust communication protocols, or causal alignment frameworks.
Figure 13 complements this diagnostic framework by synthesizing ten forward-looking design strategies aimed at mitigating these limitations. These include Retrieval-Augmented Generation (RAG), tool-based reasoning [129, 126, 127], agentic feedback loops (ReAct [132]), role-based multi-agent orchestration, memory architectures, causal modeling, and governance-aware design. Together, these three panels offer a consolidated roadmap for addressing current pitfalls and accelerating the development of safe, scalable, and context-aware autonomous systems.
Figure 12: Illustration of Chellenges: (a) Key limitations of AI Agents including causality deficits and shallow reasoning. (b) Amplified coordination and stability challenges in Agentic AI systems.
@@ V-1 Challenges and Limitations of AI Agents
While AI Agents have garnered considerable attention for their ability to automate structured tasks using LLMs and tool-use interfaces, the literature highlights significant theoretical and practical limitations that inhibit their reliability, generalization, and long-term autonomy [132, 158]. These challenges arise from both the architectural dependence on static, pretrained models and the difficulty of instilling agentic qualities such as causal reasoning, planning, and robust adaptation. The key challenges and limitations (Figure 12a) of AI Agents are as summarized into following five points:
@@
    Lack of Causal Understanding: One of the most foundational challenges lies in the agents’ inability to reason causally [172, 173]. Current LLMs, which form the cognitive core of most AI Agents, excel at identifying statistical correlations within training data. However, as noted in recent research from DeepMind and conceptual analyses by TrueTheta, they fundamentally lack the capacity for causal modeling distinguishing between mere association and cause-effect relationships [174, 175, 176]. For instance, while an LLM-powered agent might learn that visiting a hospital often co-occurs with illness, it cannot infer whether the illness causes the visit or vice versa, nor can it simulate interventions or hypothetical changes.
    This deficit becomes particularly problematic under distributional shifts, where real-world conditions differ from the training regime [177, 178]. Without such grounding, agents remain brittle, failing in novel or high-stakes scenarios. For example, a navigation agent that excels in urban driving may misbehave in snow or construction zones if it lacks an internal causal model of road traction or spatial occlusion.
@@
    Inherited Limitations from LLMs: AI Agents, particularly those powered by LLMs, inherit a number of intrinsic limitations that impact their reliability, adaptability, and overall trustworthiness in practical deployments [179, 180, 181]. One of the most prominent issues is the tendency to produce hallucinations plausible but factually incorrect outputs. In high-stakes domains such as legal consultation or scientific research, these hallucinations can lead to severe misjudgments and erode user trust [182, 183]. Compounding this is the well-documented prompt sensitivity of LLMs, where even minor variations in phrasing can lead to divergent behaviors. This brittleness hampers reproducibility, necessitating meticulous manual prompt engineering and often requiring domain-specific tuning to maintain consistency across interactions [184].
@@
    Another key limitation lies in computational cost and latency. Each cycle of agentic decision-making particularly in planning or tool-calling may require several LLM invocations. This not only increases runtime latency but also scales resource consumption, creating practical bottlenecks in real-world deployments and cloud-based inference systems. Furthermore, LLMs have a static knowledge cutoff and cannot dynamically integrate new information unless explicitly augmented via retrieval or tool plugins. They also reproduce the biases of their training datasets, which can manifest as culturally insensitive or skewed responses [186, 187]. Without rigorous auditing and mitigation strategies, these issues pose serious ethical and operational risks, particularly when agents are deployed in sensitive or user-facing contexts.
@@
    Incomplete Agentic Properties: A major limitation of current AI Agents is their inability to fully satisfy the canonical agentic properties defined in foundational literature, such as autonomy, proactivity, reactivity, and social ability [181, 142]. While many systems marketed as ”agents” leverage LLMs to perform useful tasks, they often fall short of these fundamental criteria in practice. Autonomy, for instance, is typically partial at best. Although agents can execute tasks with minimal oversight once initialized, they remain heavily reliant on external scaffolding such as human-defined prompts, planning heuristics, or feedback loops to function effectively [188]. Self-initiated task generation, self-monitoring, or autonomous error correction are rare or absent, limiting their capacity for true independence.
    Proactivity is similarly underdeveloped. Most AI Agents require explicit user instruction to act and lack the capacity to formulate or reprioritize goals dynamically based on contextual shifts or evolving objectives [189]. As a result, they behave reactively rather than strategically, constrained by the static nature of their initialization. Reactivity itself is constrained by architectural bottlenecks. Agents do respond to environmental or user input, but response latency caused by repeated LLM inference calls [190, 191], coupled with narrow contextual memory windows [161, 192], inhibits real-time adaptability. Perhaps the most underexplored capability is social ability. True agentic systems should communicate and coordinate with humans or other agents over extended interactions, resolving ambiguity, negotiating tasks, and adapting to social norms.
    However, existing implementations exhibit brittle, template-based dialogue that lacks long-term memory integration or nuanced conversational context. Agent-to-agent interaction is often hardcoded or limited to scripted exchanges, hindering collaborative execution and emergent behavior [193, 101]. Collectively, these deficiencies reveal that while AI Agents demonstrate functional intelligence, they remain far from meeting the formal benchmarks of intelligent, interactive, and adaptive agents. Bridging this gap is essential for advancing toward more autonomous, socially capable AI systems.
@@
    Limited Long-Horizon Planning and Recovery: A persistent limitation of current AI Agents lies in their inability to perform robust long-horizon planning, especially in complex, multi-stage tasks. This constraint stems from their foundational reliance on stateless prompt-response paradigms, where each decision is made without an intrinsic memory of prior reasoning steps unless externally managed. Although augmentations such as the ReAct framework [132] or Tree-of-Thoughts [160] introduce pseudo-recursive reasoning, they remain fundamentally heuristic and lack true internal models of time, causality, or state evolution. Consequently, agents often falter in tasks requiring extended temporal consistency or contingency planning. For example, in domains such as clinical triage or financial portfolio management, where decisions depend on prior context and dynamically unfolding outcomes, agents may exhibit repetitive behaviors such as endlessly querying tools or fail to adapt when sub-tasks fail or return ambiguous results. The absence of systematic recovery mechanisms or error detection leads to brittle workflows and error propagation. This shortfall severely limits agent deployment in mission-critical environments where reliability, fault tolerance, and sequential coherence are essential.
@@
    Reliability and Safety Concerns: AI Agents are not yet safe or verifiable enough for deployment in critical infrastructure [194]. The absence of causal reasoning leads to unpredictable behavior under distributional shift [195, 173]. Furthermore, evaluating the correctness of an agent’s plan especially when the agent fabricates intermediate steps or rationales remains an unsolved problem in interpretability [196, 110]. Safety guarantees, such as formal verification, are not yet available for open-ended, LLM-powered agents. While AI Agents represent a major step beyond static generative models, their limitations in causal reasoning, adaptability, robustness, and planning restrict their deployment in high-stakes or dynamic environments. Most current systems rely on heuristic wrappers and brittle prompt engineering rather than grounded agentic cognition. Bridging this gap will require future systems to integrate causal models, dynamic memory, and verifiable reasoning mechanisms. These limitations also set the stage for the emergence of Agentic AI systems, which attempt to address these bottlenecks through multi-agent collaboration, orchestration layers, and persistent system-level context.
@@ V-2 Challenges and Limitations of Agentic AI
Agentic AI systems represent a paradigm shift from isolated AI agents to collaborative, multi-agent ecosystems capable of decomposing and executing complex goals [14]. These systems typically consist of orchestrated or communicating agents that interact via tools, APIs, and shared environments [39, 18]. While this architectural evolution enables more ambitious automation, it introduces a range of amplified and novel challenges that compound existing limitations of individual LLM-based agents. The current challenges and limitations of Agentic AI are as follows:
@@
    Amplified Causality Challenges: One of the most critical limitations in Agentic AI systems is the magnification of causality deficits already observed in single-agent architectures. Unlike traditional AI Agents that operate in relatively isolated environments, Agentic AI systems involve complex inter-agent dynamics, where each agent’s action can influence the decision space of others. Without a robust capacity for modeling cause-effect relationships, these systems struggle to coordinate effectively and adapt to unforeseen environmental shifts.
    A key manifestation of this challenge is inter-agent distributional shift, where the behavior of one agent alters the operational context for others. In the absence of causal reasoning, agents are unable to anticipate the downstream impact of their outputs, resulting in coordination breakdowns or redundant computations [197]. Furthermore, these systems are particularly vulnerable to error cascades: a faulty or hallucinated output from one agent can propagate through the system, compounding inaccuracies and corrupting subsequent decisions. For example, if a verification agent erroneously validates false information, downstream agents such as summarizers or decision-makers may unknowingly build upon that misinformation, compromising the integrity of the entire system. This fragility underscores the urgent need for integrating causal inference and intervention modeling into the design of multi-agent workflows, especially in high-stakes or dynamic environments where systemic robustness is essential.
@@
    Communication and Coordination Bottlenecks: A fundamental challenge in Agentic AI lies in achieving efficient communication and coordination across multiple autonomous agents. Unlike single-agent systems, Agentic AI involves distributed agents that must collectively pursue a shared objective necessitating precise alignment, synchronized execution, and robust communication protocols. However, current implementations fall short in these aspects. One major issue is goal alignment and shared context, where agents often lack a unified semantic understanding of overarching objectives. This hampers sub-task decomposition, dependency management, and progress monitoring, especially in dynamic environments requiring causal awareness and temporal coherence.
@@
    Emergent Behavior and Predictability: One of the most critical limitations of Agentic AI lies in managing emergent behavior complex system-level phenomena that arise from the interactions of autonomous agents. While such emergence can potentially yield adaptive and innovative solutions, it also introduces significant unpredictability and safety risks [198, 153]. A key concern is the generation of unintended outcomes, where agent interactions result in behaviors that were not explicitly programmed or foreseen by system designers. These behaviors may diverge from task objectives, generate misleading outputs, or even enact harmful actions particularly in high-stakes domains like healthcare, finance, or critical infrastructure.
    As the number of agents and the complexity of their interactions grow, so too does the likelihood of system instability. This includes phenomena such as infinite planning loops, action deadlocks, and contradictory behaviors emerging from asynchronous or misaligned agent decisions. Without centralized arbitration mechanisms, conflict resolution protocols, or fallback strategies, these instabilities compound over time, making the system fragile and unreliable. The stochasticity and opacity of large language model-based agents further exacerbate this issue, as their internal decision logic is not easily interpretable or verifiable. Consequently, ensuring the predictability and controllability of emergent behavior remains a central challenge in designing safe and scalable Agentic AI systems.
@@
    Scalability and Debugging Complexity: As Agentic AI systems scale in both the number of agents and the diversity of specialized roles, maintaining system reliability and interpretability becomes increasingly complex [199, 200]. A central limitation stems from the black-box chains of reasoning characteristic of LLM-based agents. Each agent may process inputs through opaque internal logic, invoke external tools, and communicate with other agents all of which occur through multiple layers of prompt engineering, reasoning heuristics, and dynamic context handling. Tracing the root cause of a failure thus requires unwinding nested sequences of agent interactions, tool invocations, and memory updates, making debugging non-trivial and time-consuming.
    Another significant constraint is the system’s non-compositionality. Unlike traditional modular systems, where adding components can enhance overall functionality, introducing additional agents in an Agentic AI architecture often increases cognitive load, noise, and coordination overhead. Poorly orchestrated agent networks can result in redundant computation, contradictory decisions, or degraded task performance. Without robust frameworks for agent role definition, communication standards, and hierarchical planning, the scalability of Agentic AI does not necessarily translate into greater intelligence or robustness. These limitations highlight the need for systematic architectural controls and traceability tools to support the development of reliable, large-scale agentic ecosystems.
@@
    Trust, Explainability, and Verification: Agentic AI systems pose heightened challenges in explainability and verifiability due to their distributed, multi-agent architecture. While interpreting the behavior of a single LLM-powered agent is already non-trivial, this complexity is multiplied when multiple agents interact asynchronously through loosely defined communication protocols. Each agent may possess its own memory, task objective, and reasoning path, resulting in compounded opacity where tracing the causal chain of a final decision or failure becomes exceedingly difficult. The lack of shared, transparent logs or interpretable reasoning paths across agents makes it nearly impossible to determine why a particular sequence of actions occurred or which agent initiated a misstep.
    Compounding this opacity is the absence of formal verification tools tailored for Agentic AI. Unlike traditional software systems, where model checking and formal proofs offer bounded guarantees, there exists no widely adopted methodology to verify that a multi-agent LLM system will perform reliably across all input distributions or operational contexts. This lack of verifiability presents a significant barrier to adoption in safety-critical domains such as autonomous vehicles, finance, and healthcare, where explainability and assurance are non-negotiable. To advance Agentic AI safely, future research must address the foundational gaps in causal traceability, agent accountability, and formal safety guarantees.
@@
    Moreover, inter-agent dynamics themselves are susceptible to exploitation. Attackers can induce race conditions, deadlocks, or resource exhaustion by manipulating the coordination logic between agents. Without rigorous authentication, access control, and sandboxing mechanisms, malicious agents or corrupted tool responses can derail multi-agent workflows or cause erroneous escalation in task pipelines. These risks are exacerbated by the absence of standardized security frameworks for LLM-based multi-agent systems, leaving most current implementations defenseless against sophisticated multi-stage attacks. As Agentic AI moves toward broader adoption, especially in high-stakes environments, embedding secure-by-design principles and adversarial robustness becomes an urgent research imperative.
@@
    Ethical and Governance Challenges: The distributed and autonomous nature of Agentic AI systems introduces profound ethical and governance concerns, particularly in terms of accountability, fairness, and value alignment. In multi-agent settings, accountability gaps emerge when multiple agents interact to produce an outcome, making it difficult to assign responsibility for errors or unintended consequences. This ambiguity complicates legal liability, regulatory compliance, and user trust, especially in domains such as healthcare, finance, or defense. Furthermore, bias propagation and amplification present a unique challenge: agents individually trained on biased data may reinforce each other’s skewed decisions through interaction, leading to systemic inequities that are more pronounced than in isolated models. These emergent biases can be subtle and difficult to detect without longitudinal monitoring or audit mechanisms.
@@
    Immature Foundations and Research Gaps: Despite rapid progress and high-profile demonstrations, Agentic AI remains in a nascent research stage with unresolved foundational issues that limit its scalability, reliability, and theoretical grounding. A central concern is the lack of standard architectures. There is currently no widely accepted blueprint for how to design, monitor, or evaluate multi-agent systems built on LLMs . This architectural fragmentation makes it difficult to compare implementations, replicate experiments, or generalize findings across domains. Key aspects such as agent orchestration, memory structures, and communication protocols are often implemented ad hoc, resulting in brittle systems that lack interoperability and formal guarantees.
    Equally critical is the absence of causal foundations as scalable causal discovery and reasoning remain unsolved challenges [201]. Without the ability to represent and reason about cause-effect relationships, Agentic AI systems are inherently limited in their capacity to generalize safely beyond narrow training regimes [202, 178]. This shortfall affects their robustness under distributional shifts, their capacity for proactive intervention, and their ability to simulate counterfactuals or hypothetical plans core requirements for intelligent coordination and decision-making.
    The gap between functional demos and principled design thus underscores an urgent need for foundational research in multi-agent system theory, causal inference integration, and benchmark development. Only by addressing these deficiencies can the field progress from prototype pipelines to trustworthy, general-purpose agentic frameworks suitable for deployment in high-stakes environments.

### v4

@@ V Challenges and Limitations in AI Agents and Agentic AI
To systematically understand the theoretical and operational limitations of current intelligent systems, we present a comparative visual synthesis in Figure 12, which categorizes challenges and potential remedies across both AI Agents and Agentic AI paradigms. Figure 12a outlines the four most pressing limitations specific to AI Agents namely, lack of causal reasoning, inherited LLM constraints (e.g., hallucinations, shallow reasoning), incomplete agentic properties (e.g., autonomy, proactivity), and failures in long-horizon planning and recovery. These challenges often arise due to their reliance on stateless LLM prompts, limited memory, and heuristic reasoning loops.
Similarly, Figure 12b identifies eight critical bottlenecks unique to Agentic AI systems, such as inter-agent error cascades, coordination breakdowns, emergent instability, scalability limits, and explainability issues. These challenges stem from the complexity of orchestrating multiple agents across distributed tasks without standardized architectures, robust communication protocols, or causal alignment frameworks.
Figure 13 complements this diagnostic framework by synthesizing ten forward-looking design strategies aimed at mitigating these limitations. These include RAG , tool-based reasoning [135, 132, 133], agentic feedback loops (ReAct [138]), role-based multi-agent orchestration, memory architectures, causal modeling, and governance-aware design. Together, these mechanisms offer a consolidated roadmap for addressing current pitfalls and accelerating the development of safe, scalable, and context-aware autonomous systems.
Figure 12: Illustration of Challenges: (a) Key limitations of AI Agents including causality deficits and shallow reasoning. (b) Amplified coordination and stability challenges in Agentic AI systems.
@@ V-1 Challenges and Limitations of AI Agents
While AI Agents have garnered considerable attention for their ability to automate structured tasks using LLMs and interfaces to specific tools, the literature highlights significant theoretical and practical limitations that inhibit their reliability, generalization, and long-term autonomy [138, 164]. These challenges arise from both the architectural dependence on static, pretrained models and the difficulty of instilling agentic qualities such as causal reasoning, planning, and robust adaptation. These key challenges and limitations (Figure 12a) of AI Agents are summarized as follows:
@@
    Lack of Causal Understanding: One of the most foundational challenges lies in the agents’ inability to reason causally [182, 183, 184]. While LLMs, which form the cognitive core of most AI Agents are highly effective at detecting statistical correlations within training data, they do not truly understand cause-and-effect relationships. As highlighted by recent research from DeepMind and conceptual analyses by TrueTheta [185, 186, 187], these models often fail to distinguish between mere association and actual causation. For example, an LLM-powered agent might observe that hospital visits often occur alongside illness, but it cannot determine whether the illness caused the hospital visit or vice versa. More critically, such agents cannot perform counterfactual reasoning imagining what would happen if a certain intervention or change were made. This lack of causal modeling limits their ability to make informed decisions, evaluate the impact of hypothetical actions, or provide reliable recommendations in real-world scenarios where understanding “why” something happens is essential.
    Although reasoning-oriented LLMs have emerged such as DeepSeek R1 that follow a CoT approach to incrementally reason through problems, these models are not mathematically reliable reasoners (e.g., like an OWL reasoner). The chains of thought they produce are often linguistically persuasive, but not necessarily logically valid. In this sense, they do not replace formal reasoning systems such as Pellet, Bayesian networks, or causal inference frameworks, which are designed to handle logical consistency, ontological rigor, and probabilistic causality with far greater reliability.
    This limitation becomes particularly challenging under distributional shifts, where real-world conditions differ from the training regime [188, 189]. Without such grounding, agents remain brittle, failing in novel or high-stakes scenarios. For example, a navigation agent that excels in urban driving may misbehave in snow or construction zones if it lacks an internal causal model of road traction or spatial occlusion.
@@
    Inherited Limitations from LLMs: AI Agents, particularly those powered by LLMs, inherit a number of intrinsic limitations that impact their reliability, adaptability, and overall trustworthiness in practical deployments [190, 191, 192]. One of the most critical issues is the tendency to produce hallucinations, which are plausible but factually incorrect outputs. In high-stake domains such as legal consultation or scientific research, these hallucinations can lead to severe misjudgments and erode user trust [193, 194]. Compounding this is the well-documented prompt sensitivity of LLMs, where even minor variations in phrasing can lead to divergent behaviors. This brittleness hampers reproducibility, necessitating meticulous manual prompt engineering and often requiring domain-specific tuning to maintain consistency across interactions [195].
@@
    Another key limitation lies in computational cost and latency. Each cycle of agentic decision-making particularly in planning or tool-calling may require several LLM invocations. These iterations not only increase run-time latency but also scale resource consumption, creating practical bottlenecks in real-world deployments and cloud-based inference systems. Furthermore, LLMs have a static knowledge cutoff and cannot dynamically integrate new information unless explicitly augmented via retrieval or tool plugins. They also reproduce the biases of their training datasets, which can manifest as culturally insensitive or skewed responses [197, 198]. Without rigorous auditing and mitigation strategies, these issues pose serious ethical and operational risks, particularly when agents are deployed in sensitive contexts or interact directly with end users.
@@
    Incomplete Agentic Properties: A major limitation of current AI Agents is their inability to fully satisfy the canonical agentic properties defined in foundational literature, such as autonomy, proactivity, reactivity, and social ability [192, 148]. While many systems marketed as ”agents” leverage LLMs to perform useful tasks, they often fall short of these fundamental characteristics in practice. Autonomy, for instance, is typically partial at best. Although agents can execute tasks with minimal oversight once initialized, they remain heavily reliant on external scaffolding such as human-defined prompts, planning heuristics, or feedback loops to function effectively [199]. Self-initiated task generation, self-monitoring, or autonomous error correction are rare or absent, limiting their capacity for true independence.
    Proactivity is similarly underdeveloped. Most AI Agents require explicit user instruction to act and lack the capacity to formulate or re-prioritize goals dynamically based on changing context/environment or evolving objectives [200]. As a result, they behave reactively rather than strategically, constrained by the static nature of their initialization. Reactivity itself is constrained by architectural bottlenecks. Agents do respond to environmental or user input, but response latency caused by repeated LLM inference calls [201, 202], coupled with narrow contextual memory windows [167, 203], inhibits real-time adaptability.
    In addition to proactivity, social ability remains one of the most underexplored capabilities of AI Agents. Real-world AI agents and Agentic AI systems should be able to communicate and collaborate with humans or other agents over extended interactions, resolving ambiguity, negotiating tasks, and adapting to social norms. However, existing implementations exhibit brittle, template-based dialogue that lacks long-term memory integration or nuanced conversational context. Agent-to-agent interaction is often hardcoded or limited to scripted exchanges, hindering collaborative execution and emergent behavior [204, 107]. Collectively, these deficiencies reveal that while AI Agents demonstrate functional intelligence, they remain far from meeting the formal benchmarks of intelligent [205], interactive, and adaptive agents. Bridging this gap is essential for advancing toward more autonomous, socially capable AI systems.
@@
    Limited Long-Horizon Planning and Recovery: A persistent limitation of current AI Agents lies in their inability to perform robust long-horizon planning, especially in complex, multi-stage tasks. This constraint stems from their foundational reliance on stateless prompt-response paradigms, where each decision is made without an intrinsic memory of prior reasoning steps unless externally managed. Although augmentations such as the ReAct framework [138] or Tree-of-Thoughts [166] introduce pseudo-recursive reasoning, they remain fundamentally heuristic and lack true internal models of time, causality, or state evolution. Consequently, agents often fail in tasks requiring extended temporal consistency or contingency planning. For example, in domains such as clinical triage or financial portfolio management, where decisions depend on prior context and dynamically varying outcomes, agents may exhibit repetitive behaviors such as endlessly querying tools or fail to adapt when sub-tasks fail or return ambiguous results. The absence of systematic recovery mechanisms or error detection leads to brittle workflows and error propagation. This shortfall severely limits agent deployment in mission-critical environments where reliability, fault tolerance, and sequential coherence are essential.
@@
    Reliability and Safety Concerns: AI Agents are not yet safe or verifiable enough for deployment in handling or managing critical infrastructure [206]. The absence of causal reasoning leads to unpredictable behavior under distributional shift [207, 183]. Furthermore, evaluating the correctness of an agent’s plan especially when the agent fabricates intermediate steps or rationales remains an unsolved problem in interpretability [208, 116]. Safety guarantees, such as formal verification, are not yet available for open-ended, LLM-powered agents. While AI Agents represent a major step beyond static generative models, their limitations in causal reasoning, adaptability, robustness, and planning restrict their deployment in high-stakes or dynamic environments. Most current systems rely on heuristic wrappers and brittle prompt engineering rather than grounded agentic cognition. Bridging this gap will require future systems to integrate causal models, dynamic memory, and verifiable reasoning mechanisms. These limitations also set the stage for the emergence of Agentic AI systems, which attempt to address these bottlenecks through multi-agent collaboration, orchestration layers, and persistent system-level context. The persistent system-level context ensures that agents operate with a shared and evolving understanding of goals, environment, and prior decisions, enabling coherent coordination and sustained autonomy across complex workflows [209, 210]. This continuity is critical for reducing redundant processing and enabling long-horizon reasoning.
@@ V-2 Challenges and Limitations of Agentic AI
Agentic AI systems represent a paradigm shift from isolated AI Agents to collaborative, multi-agent ecosystems capable of decomposing and executing complex goals [21]. These systems typically consist of orchestrated or communicating agents that interact via tools, APIs, and shared environments [46, 25]. While this architectural evolution enables more ambitious automation, it introduces a range of amplified and novel challenges that compound existing limitations of individual LLM-based agents. The current challenges and limitations of Agentic AI systems are as follows:
@@
    Amplified Causality Challenges: One of the most critical limitations in Agentic AI systems is the magnification of lack of causal reasoning already observed in single-agent architectures. Unlike traditional AI Agents that operate in relatively isolated environments, Agentic AI systems involve complex inter-agent dynamics and cooperation, where each agent’s action can influence the decision space of others. Without a robust capacity for modeling cause-and-effect relationships, these systems struggle to coordinate effectively and adapt to unforeseen environmental shifts.
    A key issue cause by this limitation is inter-agent distributional shift, where the behavior of one agent alters the operational context for others. In the absence of causal reasoning, agents are unable to anticipate the downstream impact of their outputs, resulting in coordination breakdowns or redundant computations [211]. Furthermore, these systems are particularly vulnerable to error cascades: a faulty or hallucinated output from one agent can propagate through the system, compounding inaccuracies and corrupting subsequent decisions. For example, if a verification agent erroneously validates false information, downstream agents such as summarizers or decision-makers may unknowingly build upon that misinformation, compromising the integrity of the entire system. This fragility underscores the urgent need for integrating causal inference and intervention modeling into the design of multi-agent workflows, especially in high-stake or dynamic environments where systemic robustness is essential.
@@
    Communication and Coordination Bottlenecks: A fundamental challenge in Agentic AI lies in achieving efficient communication and coordination across multiple autonomous agents. Unlike single-agent systems, Agentic AI systems involve distributed agents that must collaboratively pursue a shared objective requring precise alignment, synchronized execution, and robust communication protocols. However, current implementations fall short in these aspects. One major issue is goal alignment and shared context, where agents often lack a unified semantic understanding of overarching objectives. This lack of shared semantic grounding hampers sub-task decomposition, dependency management, and progress monitoring, especially in dynamic environments requiring causal awareness and temporal coherence.
@@
    Emergent Behavior and Predictability: One of the most critical limitations of Agentic AI lies in managing emergent behaviors of complex system-level phenomena that arise from the interactions of autonomous agents. While such emergence can potentially yield adaptive and innovative solutions, it also introduces significant unpredictability and safety risks [212, 159]. A key concern is the generation of unintended outcomes, where agent interactions result in behaviors that were not explicitly programmed or foreseen by system designers. These behaviors may diverge from task objectives, generate misleading outputs, or even lead to harmful actions particularly in high-stake domains like healthcare, finance, or critical infrastructure.
    As the number of agents and the complexity of their interactions grow, so too does the likelihood of system instability. This includes phenomena such as infinite planning loops, action deadlocks, and contradictory behaviors emerging from asynchronous or misaligned agent decisions. Without centralized arbitration mechanisms, conflict resolution protocols, or fallback strategies, these instabilities compound over time, making the system fragile and unreliable. The stochasticity and lack of transparency in LLM-based agents further exacerbate this issue, as their internal decision logic is not easily interpretable or verifiable [213, 214]. Consequently, ensuring the predictability and controllability of emergent behavior remains a central challenge in designing safe and scalable Agentic AI systems.
@@
    Scalability and Debugging Complexity: As Agentic AI systems scale in both the number of agents and the diversity of specialized roles, maintaining system reliability and interpretability becomes increasingly complex [215, 216]. This limitation stems from the black-box chains of reasoning characteristic of LLM-based agents. Each agent may process inputs through opaque internal logic, invoke external tools, and communicate with other agents all of which occur through multiple layers of prompt engineering, reasoning heuristics, and dynamic context handling. Tracing the root cause of a failure thus requires unwinding nested sequences of agent interactions, tool invocations, and memory updates, making debugging non-trivial and time-consuming.
    Another significant constraint is the system’s non-compositionality. Unlike traditional modular systems, where adding components can enhance overall functionality, introducing additional agents in an Agentic AI architecture often increases cognitive load, noise, and coordination overhead. Poorly orchestrated agent networks where coordination, task delegation, and communication protocols are inadequately designed can result in redundant computation, contradictory decisions, or degraded task performance. Without robust frameworks for agent role definition, communication standards, and hierarchical planning, the scaling Agentic AI does not necessarily translate into greater intelligence or robustness. These limitations highlight the need for systematic architectural controls and traceability tools to support the development of reliable, large-scale agentic ecosystems.
@@
    Trust, Explainability, and Verification: Agentic AI systems pose huge challenges in explainability and verifiability due to their distributed, multi-agent architecture. While interpreting the behavior of a single LLM-powered agent is already non-trivial, this complexity is multiplied when multiple agents interact asynchronously through loosely defined communication protocols. Each agent may possess its own memory, task objective, and reasoning path, resulting in compounded opacity where tracing the causal chain of a final decision or failure becomes exceedingly difficult. The lack of shared, transparent logs or interpretable reasoning paths across agents makes it highly difficult, if not impossible, to determine why a particular sequence of actions occurred or which agent initiated a misstep.
    Compounding this opacity is the absence of formal verification tools tailored for Agentic AI. Unlike traditional software systems, where model checking and formal proofs offer bounded guarantees, there exists no widely adopted methodology to verify that a multi-agent LLM system comprising multiple large language model agents collaborating on tasks will perform reliably across all input distributions or operational contexts. This lack of verifiability presents a significant barrier to adoption in safety-critical domains such as autonomous vehicles, finance, and healthcare, where explainability and assurance are crucial. To advance Agentic AI safely, future research must address the fundamental gaps in causal traceability, agent accountability, and formal safety guarantees.
@@
    Moreover, inter-agent dynamics themselves are susceptible to exploitation. Attackers can induce race conditions, deadlocks, or resource exhaustion by manipulating the coordination logic between agents. Without rigorous authentication, access control, and sandboxing mechanisms, malicious agents or corrupted tool responses can derail multi-agent workflows or cause erroneous escalation in task pipelines. These risks are amplified by the absence of standardized security frameworks for LLM-based multi-agent systems, leaving most current implementations defenseless against sophisticated multi-stage attacks. As Agentic AI moves toward broader adoption, especially in high-stakes environments, embedding secure-by-design principles and adversarial robustness becomes an urgent research priority.
@@
    Ethical and Governance Challenges: The distributed and autonomous nature of Agentic AI systems introduces huge ethical and governance concerns, particularly in terms of accountability, fairness, and value alignment. In multi-agent settings, accountability gaps emerge when multiple agents interact to produce an outcome, making it difficult to assign responsibility for errors or unintended consequences. This ambiguity complicates legal liability, regulatory compliance, and user trust, particularly in high-stakes domains such as autonomous vehicles, scientific research, or critical infrastructure management. Furthermore, bias propagation and amplification present a unique challenge: agents individually trained on biased data may reinforce each other’s skewed decisions through interaction, leading to systemic inequities that are more pronounced than in isolated models. These emergent biases can be subtle and difficult to detect without ongoing monitoring over time or robust auditing mechanisms.
@@
    Immature Foundations and Research Gaps: Despite rapid progress and high-profile demonstrations, research and development in Agentic AI remains is still in early stage with unresolved issues that limit its scalability, reliability, and theoretical foundation. One of the central concerns is the lack of standard architectures. There is currently no widely accepted blueprint for how to design, monitor, or evaluate multi-agent systems built on LLMs . This architectural fragmentation makes it difficult to compare implementations, replicate experiments, or generalize findings across domains. Key aspects such as agent orchestration the structured coordination and role-based task allocation among agents along with memory structures and communication protocols, are often implemented in an ad hoc manner, leading to fragile systems that lack interoperability, consistency, and formal reliability guarantees.
    Equally critical is the absence of causal foundations, as scalable causal discovery and reasoning remain unsolved challenges in current AI systems [217]. Causal discovery refers to the process of identifying underlying cause-and-effect relationships from data essential for understanding how different variables influence one another. Without the ability to represent and reason about these causal links, Agentic AI systems are inherently constrained in their ability to safely generalize beyond narrow, predefined training scenarios [218, 189]. This limitation weakens their robustness when faced with distributional shifts, reduces their effectiveness in taking proactive actions, and impairs their ability to simulate alternative outcomes or hypothetical plans capabilities that are essential for intelligent coordination, adaptive planning, and high-stakes decision-making.
    The gap between functional demos and principled design thus emphasizes an urgent need for foundational research in multi-agent system theory, causal inference integration, and benchmark development. Only by addressing these deficiencies can the field progress from prototype pipelines to trustworthy, general-purpose agentic frameworks suitable for deployment in high-stake environments.